OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−6, 4) and (2, 0).

OpenStudy (saifoo.khan):

Do you know how to find the equation?

OpenStudy (anonymous):

x − 2y = −12 y = x y = x + 6 y = x + 9